Circular No. 6307

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

GRO J1744–28 D. A. Frail, Very Large Array (VLA), National Radio Astronomy Ob-

servatory; C. Kouveliotou, Universities Space Research Association; J. van Paradijs, University of Alabama at Huntsville and University of Amsterdam; and R. Rutledge, Department of Physics, Massachusetts Institute of Technology, report the detection of a time-variable radio point source in the refined (IAUC 6291; Swank and Strohmayer, private communication) XTE error box for GRO J1744–28: “We have observed the source at 3.6 cm with the VLA for three days: Feb. 2, 4, and 6 for 0.5, 0.4, and 1.0 hr, respectively. Flux densities for these three days are < 170 ± 55, 270 ± 45, and 540 ± 30 mJy, respectively. The source variability has been monitored against a second radio source in the field that remained constant during the same intervals. The location of the variable source is α = 17^{h}44^{m}36^{s}.83 ± 0^{s}.01, δ = −28^{o}45′37′′.3 ± 0′′.3 (equinox 2000.0); galactic coordinates l = 0^{o}.03, b = +0^{o}.28. The probability of an unrelated extragalactic background source in the 2.2-arcmin^{2} error box is ∼ 1 percent. Further VLA observations are planned.”

NGC 2363 L. Drissen and Y. Dutil, University of Laval; and R. Arsenault and

K. Barton, Canada-France-Hawaii Telescope (CFHT), report: “We have obtained images and spectra of the newly-discovered bright variable star in NGC 2363 (IAUC 6294) with the CFHT (+ MOS imager/spectrograph) on Feb. 7, under subarcsecond seeing conditions. This star is still very bright (V ∼ 18; M_{v} ∼ −10.0). Its spectrum is characterized by a broad Hα line in emission (broader than the nebular lines from the nearby H II region) and narrow P-Cyg profiles of He I at 667.8, 587.6, and 492.1 nm. These spectral features, combined with the high luminosity of the star and its large photometric variability, strongly suggest that this is a luminous blue variable star in outburst.”

COMET 45P/HONDA-MRKOS-PAJDUŠÁKOVÁ Total visual magnitude estimates: Jan. 27.51 UT, 8.3 (A. Hale, Cloud-

croft, NM, 10×50 binoculars; 15′ coma); 30.19, 8.4 (K. Hornoch, Lelekovice, Czech Republic, 0.35-m reflector); Feb. 1.43, 7.1 (J. E. Bortle, Stormville, NY, 10×50 binoculars; coma diameter ∼ 12′); 2.75, 7.0 (T. Lovejoy, Jimboomba, Queensland, Australia, 15×80 binoculars; coma diameter 30′).

1996 February 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green